Please help me with this problem. Caroline needs a 3% solution. How many liters of a 1% solution should she add to 2 liters of a 32% solution to get a 3% solution?
Answer by Paul(988)

Let the solution be x
x+32(2)=3(x+2)
x+64=3x+6
-2x=-58
x=29
Hence, about 29L of solution must be added.
Paul.